# Liquid: Language Models are Scalable Multi-modal Generators

### v1

@@ 4.6 Mutual Boost of Visual Generation and Understanding
In the scaling experiments, we observe that as the model size increases, the trade-off between language tasks and visual generation tasks gradually diminishes. In this section, we further explore whether there is mutual enhancement between the two tasks as the feature spaces for visual understanding and generation become unified. We conduct three sets of experiments. In the first set, we used a combination of 10M text-only data, 10M visual generation data, and 10M visual understanding data, resulting in a total of 30M data for the pre-training phase. The visual understanding data is derived by reversing the visual generation data for the captioning task, and the loss is calculated only on the text output for visual understanding task. All data are subsets of the data introduced in Section 3. The experiment, with a $1:1:1$ data ratio of the three tasks, serves as the baseline. Building on this baseline, we separately add an additional 10M visual generation data and 10M visual understanding data, forming a total of 40M data with data ratios of $1:2:1$ and $1:1:2$, respectively, labeled as ‘Add T2I’ and ‘Add I2T’. We train three GEMMA-7B models using default parameters and evaluate their performance on visual understanding and generation tasks to observe the impact of different data additions on the performance of all tasks. For the visual understanding tasks, after the pre-training phase, we perform the same instruction tuning training and then test them on standard visual understanding benchmarks.
@@ 5 Conclusion
In this paper, we present Liquid, an efficient framework enabling language models to acquire image generation and understanding capabilities without modifying the original structure. Unlike traditional multi-modal models employing extra visual models, Liquid directly tokenize images into discrete tokens that share the same embedding space with text tokens. This leads to a total unification of images and text within the models, which stokes the potential of multi-modal learning. Utilizing various existing LLMs provides a unique advantage to Liquid, enabling it to scale up easily and display similar scaling behavior to LLMs. Notably, the inherent trade-offs of unifying dual tasks within a single space diminish as the model size expands. We demonstrates that LLMs can generate high-quality images comparable to the diffusion model, and have strong potential as multi-modal generators.

### v2

@@ 3.5 Will Understanding and Generation Tasks Mutually Improve Each Other?
We conduct three sets of experiments. In the first set, we used a combination of 10M text-only data, 10M visual generation data, and 10M visual understanding data, resulting in a total of 30M data for the pre-training phase. The visual understanding data is derived by reversing the visual generation data for the captioning task, and the loss is calculated only on the text output for visual understanding task. All data are subsets of the data introduced in Section 3. The experiment, with a $1:1:1$ data ratio of the three tasks, serves as the baseline. Building on this baseline, we separately add an additional 10M visual generation data and 10M visual understanding data, forming a total of 40M data with data ratios of $1:2:1$ and $1:1:2$, respectively, labeled as ‘Add T2I’ and ‘Add I2T’. We train three GEMMA-7B models using default parameters and evaluate their performance on visual understanding and generation tasks to observe the impact of different data additions on the performance of all tasks. For the visual understanding tasks, after the pre-training phase, we perform the same instruction tuning training and then test them on standard visual understanding benchmarks.
@@ 5 Conclusion
In this paper, we present Liquid, an efficient framework enabling language models to acquire image generation and understanding capabilities without modifying the original structure. Unlike traditional multi-modal models employing extra visual models, Liquid directly tokenize images into discrete tokens that share the same embedding space with text tokens. This leads to a total unification of images and text within the models, which stokes the potential of multi-modal learning. Utilizing various existing LLMs provides a unique advantage to Liquid, enabling it to scale up easily and display similar scaling behavior to LLMs.
Leveraging this convenience, we conducted extensive scaling experiments on models ranging from 0.5B to 32B across different model families. We identified some key characteristics of multimodal models under this unified token space. 1) Firstly, by directly training LLMs on visual generation tasks, they can retain foundational language capabilities while achieving results comparable to some mainstream diffusion models. 2) Secondly, this unification of multimodal tasks tends to impair both visual generation and language tasks; however, this impairment gradually diminishes as the model size increases. 3) Lastly, we found that when visual and language tokens are represented uniformly, visual understanding and generation tasks can mutually enhance each other. This reciprocity encourages the vast potential of large-scale pretraining under this paradigm.